### v1

# Surveying the landscape of optically addressable spin qubits for quantum information and sensing technology

Abstract
Quantum technologies offer ways to solve certain tasks more quickly, efficiently, and with greater sensitivity than their classical counterparts. Yet substantial challenges remain in the construction of sufficiently error-free and scaleable quantum platforms that are needed to unlock any real benefits to society. Acknowledging that this hardware can take vastly different forms, our review here focuses on so-called spintronic (i.e. spin-electronic) materials that use electronic or nuclear spins to embody qubits. Towards helping the reader to spot trends and pick winners, we have surveyed the various families of optically addressable spin qubits and attempted to benchmark and identify the most promising ones in each group. We reveal further trends that demonstrate how qubit lifetimes depend on the material’s synthesis, the concentration/distribution of its embedded qubits, and the experimental conditions.
Keywords electron spin qubits quantum sensing spintronics materials engineering
1 Introduction
Over the last 40 years, quantum computation has progressed from concept to hardware. In 2023, IBM reported the commissioning of a 1121-qubit superconducting quantum processor[1], while Honeywell demonstrated a quantum charge-coupled device architecture based on trapped-ions[2]. Despite these achievements, the scalability and thus ultimate utility of each species of hardware remains highly contested.
Recently, spintronic (spin electronic) materials have garnered attention as contenders for qubit media. The best of these materials offer coherence times exceeding milliseconds, albeit at cryogenic temperatures[3]. Progress in the field has made the dream of usefully storing quantum information and/or implementing quantum processing at scale with appropriately interacting spins more tangible. Start-up company Quantum Brilliance[4], recently demonstrated a room-temperature (RT) qubit system based on nitrogen-vacancy (NV) centres in diamond, for example. Beyond computation, these same spintronic materials are already providing advantageous forms of sensing, especially under ambient conditions[5]. This literature review focuses, from a materials science perspective, on optimising DiVincenzo’s criterion, namely achieving long effective qubit decoherence in solid-state spin-based systems.
Acknowledging that multiple levels (qudits) can be used advantageously for error correction, quantum information is most simply encoded as the state of a two-level system. When necessary, a magnetic or electric field is applied to split otherwise degenerate sublevels. The higher and lower energy sublevels, and , form two orthogonal basis states. A pure state is formed as a coherent superposition of the two (), and the loss of this coherence results in information loss (and thus errors). For convenience, we visualise pure states as points on the surface of a Bloch sphere (Figure 2), each located by a polar angle and an azimuthal angle [6]. The value of determines the probability amplitude of finding the qubit in either the or state, while determines the relative phase of the qubit between the and states. Our ability to manipulate and exploit the quantum state of a material is ultimately limited by the lifetime of any given state, which measured by, depending on the precise application, three different time parameters, namely, the spin-lattice relaxation time, , the spin coherence time, , (also known as phase memory time, ), and the spin-dephasing time (or free induction decay time), .
1.1 Important Quantum Spin Parameters
represents the time it takes for the "longitudinal magnitude" of the state vector to decay by a factor of e. is the time required for the transverse component of the system’s spin state vector to decay by a factor of e. Put simply by DiVincenzo[7], decoherence characterises the interactions of a qubit with its environment, causing a qubit in a generic (pure) state to relax to a mixed state. This causes the loss of information in a quantum system via the loss of phase coherence. is related to the homogeneous emission linewidth, f, by 1/. Finally, is a measure of the spin coherence time that includes the effects of inhomogeneous broadening due to fluctuations in local magnetic fields, B, within spin material. According to Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio[8].
The methods used to determine , , and encompass CW and pulsed EPR[9], optically-detected magnetic resonance (ODMR)[10], and electrically detected magnetic resonance (EDMR)[11, 12]. A pulse sequence can be considered a filter that partially eliminates noise from the surroundings of a qubit system (Figure 2). Inversion and saturation recovery are common pulse sequences for measurements. Saturation recovery involves applying a strong and long pulse or a series of short pulses to achieve the equipartition of excited and ground states. Because this pulse sequence is long, the effects of spectral diffusion average out, hence the measurement closer tot the "intrinsic" can be achieved. On the other hand, inversion recovery only uses a short pulse to "flip" spins. This results in a measurement affected by spectral diffusion, hence a shorter [13]. Referring to Figure 2, the measurement sequence of only applies pulses to transform the state onto the transverse plane (-plane) and back to the longitudinal axis (z-axis). This allows the dephasing of a qubit to evolve without correction.
In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [14]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, in general, .
2 Overview of Candidate Spin Qubit Materials
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available "fully optically addressable materials" (FOAMs) that are most directly relevant to quantum information and sensing technology. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials wise, that have shown promise and report the properties of the best-performing representatives of each approach.
FOAMs are an attractive option for quantum applications since they typically involve spin-levels that are energetically separated by far more than kBT and can be initialised using visible/near-infrared light to generate quasi-"pure" quantum states. Furthermore, optical signals, being themselves far larger than kBT, are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at RT. Since the energies of the photons involved in both processes greatly exceed kBT, room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields, , is proportional to , where is duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by their photoluminescence yields or relatively short quantum spin properties. For qubit media, Weber2010 et al., developed criterion for optically addressable solid-state spin defects[15], denoting that:
1.
A state must be paramagnetic and support two or more energy levels,
2.
An optical pumping cycle can be used to initialise the qubit,
3.
Luminescence to or from the qubit state varies by qubit sublevel in some differentiable way (i.e. intensity, wavelength, or other properties),
4.
Optical transitions must not interfere with the electronic state of the host,
5.
Differences between qubit sublevels must be large enough to avoid thermal excitation.
A well-established example of a FOAM which satisfies these criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation, (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 3a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defects C3V symmetry. The application of a magnetic field lifts this degeneracy which enables the manipulation of spin density by microwave pulses thereby modify the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following futher optical excitations.
3D spin-defect FOAMs
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at RT. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electrons spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at RT[16, 17]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like" process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [18] or else second [17] order) scaling like , have been repeatedly investigated. The material’s quantum spin properties are highly dependent on the NV-centre depth[19], concentration[20], crystal strain[21], and the presence of impurities such as , N-centres or EPR-inactive neutral or positively charged NV-centres[22, 23]. Its popularity has seen a plethora of investigations to understand and modulate its spin properties. For example, high-field EPR spectroscopy and temperature-dependent measurements have demonstrated that decoherence from the N and 13C flip-flop fluctuations can be almost eliminated at low temperatures where the spin-bath is polarised[24]. Here, reached 250 s at 2 K following a sharp increase below 12 K in HTHP diamond samples. Achieving similar properties under low-field and higher temperature conditions with thermally polarised nuclei is challenging and requires meticulous materials preparation and/or the use of dynamical decoupling methods (vide infra). For example, the impact of parasitic nuclear spins and impurities was shown most remarkably by Balasuramanian et al.,[25]. Careful growth by CVD on a diamond substrate using isotopically enriched feedstock led to just 0.3 13C abundance and low levels of other paramagnetic impurity resulting in 1.8 ms at RT.
More recently, increasing attention has been paid to using diamonds as a host of other spin defects owing to its wide band gap, efficient thermal dissipation, and physical and chemical stability. As such, various magnetically-active colour centres have been investigated for quantum applications[26]. These include HV[27], BV[28], OV[29], so-called group-IV vacancy defects[30] such as SiV[31, 32], SnV[33, 34], GeV[35], and PbV[36, 37], and even transition metal defects originating as impurities like NiV[38]. To our knowledge, spin polarisation has not been observed in HV or OV centres, whilst ODMR experiments on BV and PbV have not been reported.
These metal spin-centres typically exhibit strong spin-orbit coupling (several hundred GHz) that gives rise to zero-field splitting (ZFS) even for S=1/2 species, as well as electron-nuclear spin coupling making them qudit candidates. Moreover, group-IV defects otherwise exhibit useful photonic properties that make them attractive for quantum applications such as Fourier-limited ZPL linewidths, high spectral stability, coherent photon emission, and strain-responsive band gap engineering[39]. For example, negatively charged SiV (S=1/2 system) has enabled direct observation of photon interference[40]. Outside of a dilution refrigerator, the spin coherence lifetimes are severely impacted by thermal acoustic phonon coupling. At 100 mK, reaches 1 second while the longest measured was 1.5 s. Spin coherence could be maintained by dynamical decoupling up to 600 mK where measured 60 s. can also be improved through strain engineering which modifies the spin-orbit coupling ground state splitting and subsequently the spin-phonon coupling[41]. By comparison, the neutral SiV (S=1 ground state) demonstrates an impressive 25 s, 0.1 ms even at 15 K, decreasing to 7.8 and 2 s at RT [42]. However, a route to reliably synthesising SiVs in diamonds remains elusive. Negatively charged GeV (S=1/2 ground state) has been investigated using ODMR for quantum memory applications at 300 mK and found to exhibit a 440 s and 1.46 s [43]. These defects were found to be particularly responsive to dynamical decoupling protocols with reaching 24 ms, representing a factor of two improvement compared to negatively charged SiV. SnV (S=1/2 ground state) are also robust spin centres and have been the subject of spin control experiments. Rosenthal et al., report 20 ms and 170 s when measured at 1.7 K in highly strained SnV-centres, which comprises spin lifetime for significant improvement in operation fidelity[44]. Trusheim et al., report a longer s at 1.26 ms and 540 ns at 2.9 K in a less strained system[45]. Negatively charged NiV-centres (S=1/2) are near-infrared emitters, making them especially interesting for quantum communications due to their compatibility with conventional optical cables[26]. The S=1/2 ground state of the negatively charged NiV-centre has a predicted 0.1 ms coherence time at 4 K[46], but only recently have steady-state ODMR studies been reported[38].
Beyond diamond, silicon carbide (SiC) spin systems also show promising optical and coherence properties for quantum applications[47, 48, 49, 50, 51, 52]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify a cubic and hexagonal structure, respectively, and the preceding number designates its polytype[48]. Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[53]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or V) and divacancies (VV) but can also include carbon anti-site vacancies (CVC), charged NVs, Cr4+-,[54, 55] V3+, V4+[56], Mo5+[57], and Ti-centres[58, 48, 52]. Further layers of complexity are added by consideration of additional oxidation state and crystallographic sites within each polytype. Understandably, studies often focus on the "best performing" defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the materials most significant practical limitation as a spin qubit candidate compared to other FOAM platforms.
Nevertheless, several defects exhibit outstanding spin-optical properties and the fabrication challenges are beginning to be addressed. Negatively charged V in isotopically purified (28Si) 4H-SiC exhibit the largest ODMR contrast at 97 at 4 K[53]. With a S=3/2 ground state, this defect exhibits ZFS associated of a few MHz with degenerate pairs of <1/2| and <3/2| states. Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called "V1" state, <3/2| shift higher energy than the <1/2| states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes <3/2| states to selectively decay by into a non-radiative metastable state followed by spin-selective repopulation of <1/2| ground states thereby generating a spin polarisation of up to 90. The authors record and of 0.80.12 ms and 302 s, respectively. More recently, it was shown that V defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties[59]. Here, of bulk V-centres was measured at 344 s at 10 K, whilst those in the 1 s diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at RT[60].
By comparison, divacancies also exhibit compelling spin-optical properties albeit with lower optical contrasts. Divacancies are formed by annealling pre-irradiated SiC at over 700oC[61, 62]. Conversion efficiency into divacancies only reaches a few percent, which may in part be due to counterproductive divacancy-dissociation back into VC, V, and CVC-centres during the annealing process[63]. As S=1 ground state species, divacancies exhibit more intense luminescence from their mS= states compared to their mS=0 state and can optically initialised the same fashion as NV- diamond thanks to its C-symmetry. The kk-divacancy (a neutral VV-centre) in 4H-SiC with natural isotopic abundance can demonstrate times of 1.3 ms at 20 K when decoupled from 13C and 29Si nuclear spins under a 30 mT field[64], significantly longer than NV- diamond under similar conditions. There are also several unknown undefined divacancy-type centres that demonstrate remarkable insensitivity to temperature. For example, Yan et al., reported that PL8-centres, which are triplet ground state species, exhibit similar quantum spin properties at 20 K, with measured and of 15.60.5 s and 18410 ns, and at RT, with measured and of 9.10.1 s and 1809 s[65]. This long-lived RT spin coherence is shared by PL6-centres which are distinguishable by their 1038 nm (vs PL8’s 1007 nm) ZPL emission, demonstrating that several species in a SiC matrix are suitable for RT quantum applications. Moreover, divacancy systems were the first spin-centres demonstrated to be amendable to all electrical spin-ensemble readout and initialisation[66, 67]. These functionalities have since been discovered with monovacancies under ambient conditions[68], demonstrating a strong potential to avoid some of the difficulties associated with ODMR spectroscopy such as photon collection efficiency.
Beyond vacancy systems, several metal ion centres have demonstrated interesting spin-optical properties. By far the best however is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits >1 second, and and = 81 s and 317 ns, respectively, at 15 K[55]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath[54].
p- and d-block Molecular FOAMs
Molecular systems are becoming increasingly popular and offer an enticing opportunity to develop chemically tuneable quantum materials catered to different applications[98, 99, 100, 101]. Ground-up synthesis enables the incorporation of particular functionalities such as stable radicals[102, 103], modulation of triplet/singlet yields, enrichment with low or zero nuclear magnetic moments such as deuterium, oxygen and sulfur, or the targeted inclusion of nuclear spin-active elements such as nitrogen, phosphorous and several transition metals and lanthanides (vide infra). A synthetic approach also enables changes to the host matrix[73], spin concentration, defect orientation, and material processing approaches which are limited for defect-based systems. For example, single-molecule ODMR spectroscopy of pentacene molecules in a p-terphenyl matrix (Pc:PTP) has already been demonstrated at cryogenic temperature in a series of remarkable works by Wratchup and colleagues[104, 105, 106]. Only recently have pulsed experiments been performed to reveal highly competitive contrast and spin coherence properties at RT. The Tx-Ty spin transition of a 0.1 crystal Pc:PTP demonstrates 23 s, 2.7 s and 500 ns by pulsed ODMR spectroscopy[69, 107]. Investigations using the more strongly spin polarised Tx-Tz transition in both crystals and 100 nm-thin films at 0.01 and 0.1, respectively, reveal similar spin dynamics and also suggest an ability to modulate Pc:PTP’s spin properties according to sample thickness and spin concentration[70].
Further organic systems have demonstrated potential as FOAMs, though to our knowledge pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix[108]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene[109] and dinaphtho-(2’,3’:1,2);(2",3":6,7)-pyrene[110]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright RT ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics[111, 112]. These materials exhibit narrow resonance lines due to intermolecular site hopping. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured[111]. Their high spin densities (50) are highly advantageous for quantum sensing where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for exotic spin behaviours like singlet fission and triplet-triplet annihilation[113]. For example, at RT the triplet states of Phenazine:TCNQ are predominately formed by singlet fission. Despite high spin densities, experiments using transient nutation ESR spectroscopy have been used to estimate and times of 1 s and 600 ns, respectively[114]. More recently, several ground state radical and diradical materials with an optical readout capacity have also been demonstrated[115, 71], as have materials with quintet states[116, 117, 118].
FOAMs consisting of d-block have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state[119]. To reduce dipole coupling, the spin-active moiety is diluted in a matrix of the S=0 isostructural tin analogue with crystals grown from hexane solutions[72]. Optical spin polarisation is then achieved by exciting molecular spins between the S=1 ground state and S=0 excited state, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins which benefit from an E=0 triplet state and a relatively long of 0.22 ms at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast, enabling the measurement of of 640 ns. The coherent properties can be significantly improved by using a non-isostructural tin(4-fluoro-2-methylphenyl)4 host matrix[120]. Here, the ZFS is significantly increased giving rise to clock transitions similar to those realised by the ZEFOZ method. As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively.
Recently, even fluorescent proteins have been demonstrated as a viable spin qubit media in vivo and in solution at room temperature. Feder et al., used ODMR spectroscopy to demonstrate that at 80 K the X-Z and Y-Z triplet transitions of enhanced yellow fluorescent protein (EYFP) demonstrated a 44 and 32 ODMR contrast, respectively[97]. This corresponds with zero-applied field (estimated from spin polarisation decay) and of 141 and 1.5 s. Using Carr-Purcell-Meiboom-Gill (CPMG) dynamical decoupling the effective decoherence time () reached 16 s. The authors largely circumvent overhead limitations from long triplet lifetimes (ms) using an additional near-infrared pulse to induce T1 to T2 transitions thereby inducing reverse intersystem crossing and delayed fluorescence from S1. This interesting methodology could be suitable for other spin systems with quasi-resonant S1 and T2 electronic states, such as pentacene, to increase their sensitivity by increasing measurement repetition rates.
f-block molecular FOAMs
Lastly, there has been significant interest in d- and f-block molecular systems. Here, we make a distinction between molecular systems, where the ion is doped into a molecular lattice, and trapped ion systems which can give rise to extremely long coherence times, but use radically different device architectures and are not subject to qubit-quality improvements through molecular engineering[121, 122, 123]. Like organic materials, d- and f-block molecular systems benefit from diverse chemistry and processing methods that can in principle reduce decoherence, enable built-in error correction, and lead to scalable qubit media. Moreover, these heavy elements often exhibit strong spin-orbit coupling leading to high magnetic anisotropy that can protect spin states from small magnetic fluctuations[124, 125]. However, as a result of these effects and unlike the candidates previously discussed, the optical transitions in question often correspond to nuclear spin transitions. Examples of optically active materials benefit from narrow and stable spin-dependent emission profiles at near-infrared frequencies making them potentially compatible with conventional telecom optical fibres. Colour centres for which optical addressability has been established include Cr4+[72], Eu3+-[126], Pr3+-[92], Er3+-[89], Yb3+-[127], and Sm3+:Y2SiO5 [128]. These lanthanide-based centre benefit from highly shielded f-orbital electrons leading to relatively long coherence times at low temperatures[129], as well as strong hyperfine coupling (tens of MHz) with I=5/2 (Eu, Pr, 173Yb), 1/2 (171Yb), 7/2 (149Sm, 167Er) nuclei combined with strong spin-orbit coupling leading to discreet optically addressable spin-sublevels. Y2SiO5 has been favored due to the ability to grow large crystals with excellent optical properties by the Czochralski method[130]. However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath and inhomogeneity of applied-magnetic fields, various decoupling techniques have emerged[131, 132, 87]. Perhaps the most successful is the so-called zero first-order Zeeman (ZEFOZ) technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [133]. In this "clock transition" regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire >>100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is in light-transport of quantum states[134].
Interestingly, the larger magnetic moment of Pr3+ can give rise to "frozen core"-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins[131]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K[92]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K[131] and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit[135]. Coherence times can be further enhanced by using a host matrix whereby the principle host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K[93].
Spin coherence times of Er3+:Y2SiO5 have so far been shorter than the best performing Eu3+ materials with measured at 1.30.01 seconds at 1.4 K, despite exhibiting the frozen core effect[89]. However, this was achieved without using the ZEFOZ method and therefore these times can likely be significantly extended. Er3+ has also been studied in a Y2O3 host with reaching 140 s at 1.8 K and with an applied field of 4 T[88]. Decoherence was dominated by phonon-driven dipole-dipole interactions and the nuclear spin bath at high fields, similar to Er3+:KTiOPO4 where was measured at 200 s under similar conditions[90].
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure resulting in a simple 4-level system. Of the 171Yb3+-doped materials[95], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K and 90 mT, which quickly increases above 4 K due to Raman relaxation where is ultimately limited by [94]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling until reached 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extend 1 ms[127]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective "hole burning" in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 spin polarisation, thereby effectively generating mK spin temperatures. The result is an increase in the optical from 0.3 to 0.8 ms [136]. To our knowledge, pulsed optical decoherence studies have not been performed on Sm3+:Y2SiO5, however, its I=7/2 nucleus and strong hyperfine coupling may be useful for qudit systems, and it is predicted to be less sensitive to magnetic field fluctuations that Er3+[128, 137].
3 Guidelines for Optimising Performance of Spin-Qubit Materials
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and spin-bath dependence of the spin properties and the difference in fluorescence between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / since 1/ = 1/ + B (Figure 5), where a low ratio indicates that is close to . Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated van der Waals materials where atomic-scale directional anisotropy ensures that all molecules "feel" the same magnetic environment. By comparison, there are significant difficulties associated with synthesising aligned 3D-spin defects in materials such as SiC and diamond.
Further improvements in spin parameters can be obtained by positional engineering of defect centres. Clustering (or the straggling) of spin-active dopants in a substrate poses spin-spin coupling from the environment (nuclear spins), which decreases . Controlled doping becomes crucial in decreasing spin density around spin-active defect centres. Eliminating unwanted spins like nuclear spins requires isotopic purity of substrate material or the careful doping of spin active centres or defects within the host matrix. More recently, Plasma-Enhanced CVD methods have been used to achieve higher deposition rates while minimising the impact of energetic ions or electrons affecting the colour centres in NV and SnV diamonds[138]. Another novel method to precisely control the position of each defect is laser writing. Aberration-corrected optics allow for the precise positioning of vacancies in diamond systems, with a 45 success probability of a vacancy being located within 200 nm of a desired position[139]. Therefore, chemical systems with charge transfer or hydrogen bonding motifs may yet demonstrate advantages due to their ability to effectively engineer the placement of molecules in 3D. Moreover, as seen with Pc:PTP, and Pr3+:Y2SiO5 vs. the La2(WO5)3 host, materials engineers should avoid defect-site strain by selecting hosts with closely matched physical parameters to the defect. Finally, further improvements can readily be realised through dynamical decoupling methods that are tailored for each application. For example, while engineering clock transitions using magnetic fields or zero-field splitting is appealing for quantum optics where spectral stability is prized, it is not necessarily useful for sensing or information processing. This is because the associated reduced magnetic field sensitivity would reduce the quantum operation fidelity. On the other hand, focused electromagnetic driving of parasitic impurities, such as N-centres in diamond, has yet to be significantly explored in molecular systems. Significant improvements in and would likely emerge from a combination of field-driving and pulsed refocusing methods such as CPMG.
4 Conclusion and Outlook
Optically addressable electron spin systems show great potential as a diverse form of spin qubit media for quantum sensing, communications, and information processing applications. From this review, we have attempted to benchmark the different families of materials and identify useful investigative and experimental approaches that can be translated across the field. The most significant hurdles faced by chemists and materials engineers remains the strong temperature dependence of the spin-lattice relaxation and thermal polarisation of the spin bath, which renders most d- and f-block metal systems unpracticable above a few kelvin, below which these materials demonstrate the longest coherence times by a significant margin. However, above liquid helium temperatures, it is clear that light-element materials such as colour centres in diamond, SiC and most recently, molecular systems demonstrate more robust quantum spin parameters. Nevertheless, significant advancements in dynamic decoupling techniques within the last 20 years and isotope engineering have enabled the realisation of remarkably competitive quantum spin properties. Within the next decade, it should be expected that further advances and knowledge transfer between investigators of different material platforms with lead to devices capable of significantly impacting society, especially in the field of magnetic field sensors and coherent quantum optics. We hope also that by collating (meta)data on the available spin qubit candidates that the field will eventually benefit from the processing and predictive power of machine learning and artificial intelligence approaches to materials discovery.
5 Acknowledgements
This work was supported by the UK Engineering and Physical Science Research Council through grants EP/V048430/1 and EP/W027542/1. The authors are also very grateful to Dr. Daan M. Arroo for fruitful discussions and advice.