### v2

# Surveying optically addressable spin qubits for quantum information and sensing technology

Abstract
Quantum technologies offer ways to solve certain tasks more quickly, efficiently, and with greater precision than their classical counterparts. Yet substantial challenges remain in the construction of sufficiently error-free and scalable quantum platforms needed to unlock any real benefits to society. Acknowledging that this hardware can take vastly different forms, our review here focuses on so-called spintronic (i.e. spin-electronic) materials that use electronic or nuclear spins to embody qubits. Towards helping the reader to spot trends and pick winners, we have surveyed the various families of optically addressable spin qubits and attempted to benchmark and identify the most promising ones in each. We go on to reveal further trends that demonstrate how qubit lifetimes depend on the material’s synthesis, the concentration/distribution of its embedded qubits, and the experimental conditions.
Keywords spin qubits quantum sensing spintronics materials engineering
1 Introduction
Over the last 40 years, quantum computation has progressed from concept to hardware. In 2023, IBM reported the commissioning of a 1121-qubit superconducting quantum processor [1], while Honeywell demonstrated a quantum charge-coupled device architecture based on trapped ions [2]. Despite such achievements, the scalability and, thus ultimate utility of each species of hardware remains contested.
Recently, spintronic (spin-electronic) materials have garnered attention as contenders for use as qubit media. The best of these materials offer coherence times exceeding milliseconds, albeit at cryogenic temperatures [3]. Progress in the field has made the dream of usefully storing quantum information and/or implementing quantum processing at scale with appropriately interacting spins more tangible. For example, the start-up company Quantum Brilliance recently demonstrated a room-temperature qubit system based on nitrogen-vacancy (NV) centres in diamond [4]. Beyond computation, these same spintronic materials are already providing advantageous forms of quantum sensing, especially under ambient conditions where thermally-driven electronic and photonic noise present significant challenges for quantum applications [5]. Their advantages stem from the ability to optically initialise spin states into highly non-Boltzmann populations and, likewise, to implement optical read-out routines that are less impacted by blackbody radiation.
In quantum spintronic materials, information is most simply encoded in the spin state of a two-level system, but can be encoded across multiple levels in high-spin (S1) systems, so facilitating error correction [6]. For qubits, a magnetic or electric field is applied to split otherwise degenerate spin states. The higher and lower energy spin states, and , form two orthogonal basis states. A pure state is formed as a coherent superposition of the two (, see Figure 1a). Geometrically, we can visualise any such state as a point on the surface of a Bloch sphere, the location being defined by a polar angle and an azimuthal angle [7]. The value of determines the probability amplitude of finding the qubit in either the or state, while determines the relative phase of the qubit between the and states.
To be useful, spintronic materials must enable both single-spin addressability and long storage times at high levels of fidelity. These properties are, in turn, limited by the materials’ spin-lattice relaxation () and spin decoherence times (, also known as phase memory time). and represent the time it takes for the "longitudinal" and "transverse" magnitude of the spin state vector, respectively, to decay by a factor of e. In practical terms, is the lifetime of a non-Boltzmann population (Figure 1b), and, put simply by DiVincenzo [8], characterises the interactions of a qubit with its environment. These interactions are ultimately responsible for causing a qubit in a generic (pure) state to relax to a mixed state, leading to the loss of quantum information via phase decoherence (Figure 1c). These parameters can be measured using electron spin resonance (ESR) techniques, which measure light/microwave absorptions and emissions from a material as a function of frequency or magnetic field strength (Figure 2). is typically measured by Hahn-Echo pulses but can also be determined from the homogeneous emission linewidth of its resonance frequency, (Figure 1d). In practice, spin systems exhibit an inhomogeneously broadened linewidth, , due to differences in the magnetic field, B, at the locations of otherwise equivalent spins (see Figure 1e). These differences are in turn due to differences in the locations and orientations of the electronic and/or nuclear spins that surround each measured spin. Following Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio [9]. In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [10]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, .
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available "fully optically addressable materials" (FOAMs) that are most directly relevant to quantum information and sensing technology [11]. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials-wise, that have shown promise and report the properties (where available) of the best-performing representatives of each approach.
2 Overview of Candidate Spin Qubit Materials
FOAMs are an attractive option for quantum applications since they typically involve spin levels that are energetically separated by far more than and can be initialised using visible/near-infrared light to generate quasi-"pure" quantum states. Furthermore, optical signals, being themselves far larger than , are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at room temperature. Since the energies of the photons involved in both processes greatly exceed , room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields (for example), , is proportional to , where is the duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by either low photoluminescence yields or relatively short-lived spin polarizations () and coherences (), which have been summarised in Figure 3 to support our discussion. As a prerequisite for inclusion in figure 3, we have opted to follow the criteria put forward by Weber et al., for the realization of an optically addressable solid-state spin qubit (that is useable)[12]:
1.
A state must be paramagnetic and support two or more energy levels,
2.
An optical pumping cycle can be used to initialise the qubit,
3.
Luminescence to or from the qubit state varies by qubit sublevel in some differentiable way (i.e. intensity, wavelength, or other properties),
4.
Optical transitions must not interfere with the electronic state of the host,
5.
Differences between qubit sublevels must be large enough to avoid thermal excitation.
2.1 3D spin-defect FOAMs
2.1.1 Diamond spin centres
A well-established example of a FOAM which satisfies the above criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 4a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defects C3V symmetry. The application of a magnetic field lifts this degeneracy, which enables spin manipulation using microwave pulses, thereby modifying the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following subsequent optical excitations and fluorescence back to T0.
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at room temperature. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electron spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at room temperature [60, 61]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like" process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [62] or else second [61] order) scaling like , have been repeatedly investigated.
The material’s quantum spin properties are highly dependent on the NV-centre depth [63], concentration [64], crystal strain [65], and the presence of impurities such as 13C, N-centres or EPR-inactive neutral or positively charged NV-centres [66, 67]. Its popularity has seen a plethora of investigations to understand and modulate its spin properties. For example, high-field EPR spectroscopy and temperature-dependent measurements have demonstrated that decoherence from the N and 13C flip-flop fluctuations can be almost eliminated at low temperatures where the spin-bath is polarised [68]. Here, reached 250 s at 2 K, following a sharp increase below 12 K in high temperature high pressure (HTHP) diamond samples. Achieving similar properties under low-field and higher temperature conditions with thermally polarised nuclei is challenging and requires meticulous materials preparation and/or the use of dynamical decoupling methods. For example, the impact of parasitic nuclear spins and impurities was shown most remarkably by Balasuramanian et al., [22]. Careful growth by CVD on a diamond substrate using isotopically enriched feedstock led to just 0.3 13C abundance and low levels of other paramagnetic impurity, resulting in 1.8 ms at room-temperature.
More recently, increasing attention has been paid to using diamonds as a host for other spin defects owing to diamond’s wide band gap, efficient thermal dissipation, and physical and chemical stability. As such, various magnetically-active colour centres have been investigated for quantum applications [69]. These include HV [70], BV [71], OV [72], so-called group-IV vacancy defects [73] such as SiV [74, 75], SnV [76, 57], GeV [77], and PbV [78, 79], and even transition metal defects originating as impurities like NiV [80]. To our knowledge, spin polarisation has not been observed in HV or OV centres, whilst ODMR experiments on BV and PbV have not been reported.
Metal-vacancy spin-centres typically exhibit strong spin-orbit coupling (several hundred GHz) that gives rise to zero-field splitting (ZFS) even for S=1/2 species, as well as electron-nuclear spin coupling, making them qudit candidates (Figure 4b-d). Moreover, group-IV defects otherwise exhibit useful photonic properties that make them attractive for quantum applications such as Fourier-limited zero-phonon line (ZPL) linewidths, high spectral stability, coherent photon emission, and strain-responsive band gap engineering [81].
For example, negatively charged SiV (S=1/2 system, (Figure 4c) has enabled direct observation of photon interference [82]. Outside of a dilution refrigerator, the spin coherence lifetimes are severely impacted by thermal acoustic phonon coupling. At 100 mK, reaches 1 second while the longest measured was 1.5 s. Spin coherence could be maintained by dynamical decoupling up to 600 mK where the coherence time during a dynamical decoupling scheme () measured 60 s. can also be improved through strain engineering, which modifies the spin-orbit coupling, which is responsible for inducing the ground state splitting, and subsequently the spin-phonon coupling [83]. By comparison, the neutral SiV (S=1 ground state) demonstrates an impressive 25 s, 0.1 ms even at 15 K, decreasing to 7.8 and 2 s at room temperature [23]. However, a route to reliably synthesising SiVs in diamonds remains elusive.
Negatively charged GeV (S=1/2 ground state, (Figure 4d) has been investigated using ODMR for quantum memory applications at 300 mK and found to exhibit a 440 s and 1.46 s [26]. These defects were found to be particularly responsive to dynamical decoupling protocols with reaching 24 ms, representing a factor of two improvement compared to negatively charged SiV.
SnV-s (S=1/2 ground state) are also robust spin centres and have been the subject of spin control experiments. Rosenthal et al., report 20 ms and 170 s when measured at 1.7 K in highly strained SnV-centres, which comprises spin lifetime for significant improvement in operation fidelity [25]. Trusheim et al., report a longer 1.26 ms and 540 ns at 2.9 K in a less strained system [58].
Negatively charged NiV-centres (S=1/2) are near-infrared emitters, making them especially interesting for quantum communications due to their compatibility with conventional optical cables [69]. The S=1/2 ground state of the negatively charged NiV-centre has a predicted 0.1 ms coherence time at 4 K [84], but only recently have steady-state ODMR studies been reported [80].
2.1.2 Silicon Carbide Spin Centres
Beyond diamond, silicon carbide (SiC) also shows promising optical and coherence properties for quantum applications [85, 86, 87, 88, 89, 90]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify the cubic and hexagonal structures, respectively, and the preceding number designates its polytype [86] (see Figure 5a). Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[31]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or VSi) and divacancies (VV) but can also include carbon anti-site vacancies (CSiVC) (see Figure 5b and c), charged NVs, Cr4+- [41, 91], V3+, V4+ [92], Mo5+ [44], and Ti-centres [93, 86, 90]. Further layers of complexity are added by consideration of additional oxidation states and crystallographic sites within each polytype. Understandably, studies often focus on the "best performing" defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the material’s most significant practical limitation as a spin qubit candidate compared to other platforms.
Nevertheless, several defects exhibit outstanding spin-optical properties, and the fabrication challenges are beginning to be addressed. Negatively charged VSi in isotopically purified (28Si) 4H-SiC exhibits the largest ODMR contrast at 97 at 4 K [31]. With an S=3/2 ground state, its associated ZFS is a few MHz with degenerate pairs of |1/2 and |3/2 states (Figure 5d). Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called "V1" state, |3/2 shifts higher energy than the |1/2 states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes |3/2 states to selectively decay into a non-radiative metastable state, followed by spin-selective repopulation of |1/2 ground states. This leads to a spin polarisation of up to 90 which is paired with record s and s of 0.80.12 ms and 302 s, respectively. More recently, it was shown that VSi defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties [32]. Here, of bulk VSi-centres was measured at 344 s at 10 K, whilst those in the 1 m diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor of 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at room-temperature [33].
By comparison, divacancies also exhibit compelling spin-optical properties, albeit with lower optical contrasts. Divacancies are formed by annealing pre-irradiated SiC at over 700oC [97, 98]. Conversion efficiency into divacancies only reaches a few percent, which may in part be due to counterproductive divacancy-dissociation back into VC, VSi, and CSiVC-centres during the annealing process [99]. As S=1 ground state species, divacancies exhibit more intense luminescence from their mS= states compared to their mS=0 state and can be optically initialised in the same fashion as NV-diamond thanks to its C3V-symmetry. The kk-divacancy (a neutral VSiV-centre) in 4H-SiC with natural isotopic abundance can demonstrate times of 1.3 ms at 20 K when decoupled from 13C and 29Si nuclear spins under a 30 mT field [37], significantly longer than NV-diamond under similar conditions. There are also several divacancy-type centres that demonstrate remarkable insensitivity to temperature. For example, Yan et al., reported that PL8-centres, which are triplet ground state species, exhibit similar quantum spin properties at 20 K, with measured and of 15.60.5 s and 18410 ns, and at room-temperature, with measured and of 9.10.1 s and 1809 s [36] (Figure 5e-g). This long-lived room-temperature spin coherence is shared to a reduced extent by PL6-centres, which are distinguishable by their 1038 nm (vs PL8’s 1007 nm) ZPL emission, demonstrating that several species in a SiC matrix are suitable for room-temperature quantum applications. Moreover, divacancy systems were the first spin-centres demonstrated to be amenable to an all-electrical spin-ensemble readout and initialisation scheme [100, 101]. These functionalities have since been discovered with monovacancies under ambient conditions [102], demonstrating a strong potential to avoid some of the difficulties associated with ODMR spectroscopy, such as pump-light stability and photon collection efficiency.
2.1.3 Inorganic Semiconductor Spin Centres
Beyond vacancy systems, several metal-ion centres have demonstrated interesting spin-optical properties. By far the best, however, is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits >1 second, and and = 81 s and 317 ns, respectively, at 15 K [91]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath [41].
2.2 2D- and van der Waals FOAMs
Van der Waals (vdW) materials are made of 2D-atomically thin layers weakly bound together through van der Waals interactions. These include graphene, hexagonal boron nitride (hBN) and transition metal dichalcogenides (TMDs). Despite a range of remarkable materials having been identified as ODMR-active, the precise structure and even spin multiplicity of the spin species is often ambiguous. Nevertheless, like other defect spin-qubit systems, inhomogeneous broadening and spin-phonon interactions are the principal causes of spin decoherence.
The prototypical vdW material, graphene, was an early candidate for harbouring spin qubits; however, the metallic band structure of graphene is not conducive to forming luminescent spin centres. Instead, its isoelectronic analogue, hBN, has been the centre of remarkable advancement as a 2D-optically addressable platform at room temperature, building off its success as a single-photon source [103]. Magnetically-active spin centres in hBN were identified by EPR in the 1960s, with hyperfine structures giving clues to their identity as S=1/2 boron centres and carbon impurities [104, 105, 106]. However, the first optically addressable spin centres were only identified recently as triplets stemming from a negatively charged vacancy, likely originating from a missing boron atom, so-called -centres [107, 108, 109, 110] (Figure 6a). Similarly to NV-diamond centres, optical initialisation is achieved by photoexcitation with 532 nm light, leading to ISC into a metastable singlet, which spin-selectivity repopulates the fluorescent ms=0 state (TZ at zero-applied field, Figure 6b). The readout of the spin state occurs through 850 nm fluorescence with pulsed ODMR contrast of 1 at room temperature and almost 30 at 10 K, revealing a ZFS of |D|3.5GHz (Figure6c). Spin-relaxation and Hahn-echo measurements reveal a room-temperature at of 18 s and of 2 s [27]. These properties are particularly remarkable since both boron and nitrogen exhibit isotopically-abundant nuclear spins ( = 1, = 3, and = 3/2). Suppression of the nuclear spin bath using a second microwave source centred on a hyperfine line was found to improve and up to 25 s and 7.5 s, respectively. Work continues optimising the synthesis conditions for hBN to enhance the quantum properties of the defects. For instance, the contrast of -centres can be increased to 10 at room temperature by controlling the material purity and electron irradiation dose, though appears to remain phonon limited [111].
Due to hBN’s potential as a tunable single photon source, significant effort has been made to control carbon-defect structures through careful synthesis. CVD-grown hBN has been found to contain a plethora of optically-addressable carbon-based spin species with stable ZFLs in the visible range; though most have yet to be conclusively structurally characterised, or even the multiplicity conclusively determined (see Figure 6d for candidate structures and Figure 6e for the spin initialisation scheme). For example, three distinct so called ""-defects, likely stemming from the substitution of a boron or nitrogen atom by carbon ("D1", "D2", "D3", all likely S=1/2), demonstrate similar spin relaxation and coherence quantum properties (0.2 s) with a continuous wave (CW)-optical contrast up to 20 at 5 K [30] (Figure 6f). Their inhomogeneous linewidths are also limited by the nuclear spin bath (40-60 ns), where local spins also bestow an apparent zero-field splitting of between 8-10 MHz.
Though ensemble measurements remain challenging due to carbon-doped hBN’s broad emission profile, a recent report has shown that it is possible to survey over 400 defect centres in CVD-grown samples. Collectively known as ""-centres, careful measurements of individual spin-centres reveal distinctive triplet character with a ZFS of 15 MHz in contrast to most other accepted carbon-defects found in hBN [112]. The triplet identity was later unambiguously characterised and confirmed for a series of defects with significantly larger ZFS with |D|2 GHz and |E|50-80 MHz [29] (Figure 6g and h). The room-temperature was measured to be 20030 s, 1 s and 100 ns, competitive with -triplet centres. However, these properties are matched with a remarkable ODMR contrast of almost 50 (though it decreases with an applied magnetic field), giving rise to projected D.C. magnetic field sensitivity of up to 3 T Hz-1/2. A recent survey of similar defects demonstrated that under high microwave and optical pump powers, the optical contrast could reach as high as 90, though most exhibit <50 despite similar ZFS, and [113]. The authors explore the potential of this material as a medium for D.C. vectorial quantum magnetometry by taking advantage of the defects’ low symmetry and in-plane quantisation axis. Though yet-to-be experimentally verified, the authors hypothesise that applying a field bias along a different material axis corresponding to each of the three triplet states, Tx, Ty, and Tz, should result in subtle changes in contrast that can be related to particular Tn states to yield 3-vector components of the applied magnetic field.
More complex carbon species demonstrating similar coherent behaviour at room temperature have been explored using metal-organic vapour-phase epitaxy with carbon-based additives [114, 28]. The so-called ""-defect (potential identity ) is another potential S=1/2 species or, two weakly coupled S=1/2 centres, has recently been reported as being co-located with -centres following electron irradiation of hBN flakes [28]. While their and are similar to -species (13 s and 80 ns, respectively), it can be used to "sense" -centres through spin cross relaxation at an appropriate applied magnetic field. Due to its out-of-plane magnetic quantisation axis (versus the in-plane axis of -centres), it can also more effectively be used to image magnetic structure on underlying substrates with an estimated DC magnetic sensitivity of 1 T Hz-1/2.
2.3 Molecular FOAMs
2.3.1 p-block molecular spin centres
Molecular systems are becoming increasingly popular and offer an enticing opportunity to develop chemically tunable quantum materials catered to different applications [115, 116, 117, 118]. Ground-up synthesis enables the incorporation of particular functionalities such as stable radicals [119, 120], modulation of triplet/singlet yields, enrichment with low or zero nuclear magnetic moments such as deuterium, oxygen and sulfur, or the targeted inclusion of nuclear spin-active elements such as nitrogen, phosphorous, several transition metals, and lanthanides (vide infra). A synthetic approach also enables changes to the host matrix [19], spin concentration, defect orientation, and material processing approaches, which are limited for defect-based systems.
Single-molecule ODMR spectroscopy of pentacene molecules in a p-terphenyl matrix (Pc:PTP) has already been demonstrated at cryogenic temperatures in a series of remarkable works by Wrachtrup and colleagues [121, 122, 123]. Only recently have pulsed experiments been performed to reveal highly competitive contrast and spin coherence properties at room-temperature. The Tx-Ty spin transition of a 0.1 crystal Pc:PTP demonstrates 23 s, 2.7 s and 500 ns [13, 124]. Investigations using the more strongly spin polarised Tx-Tz transition in both crystals and 100 nm-thin films at 0.01 and 0.1, respectively, reveal similar spin dynamics and also suggests an ability to modulate Pc:PTP’s spin properties according to sample thickness and spin concentration [14]. These robust spin-optical properties can be maintained in Pc:PTP nanocrystals down to 200 nm with an optical contrast similar to equivalently sized NV-diamond nanoparticles [125]. When coated with a Pluronic F-127 polymer, these nanoparticles also display excellent biocompatibility with low toxicity over 72 hours according to in vitro explorations, paving the way for their use in cell assay techniques.
The potential of using synthetic chemistry to modulate the spin dynamics to improve spin-optical properties has also been explored. Substituting two carbon atoms on pentacene for nitrogen atoms resulted in 6,13-diazapentacene, which could be doped into a p-terphenyl host (DAP:PTP) by crystal growth and vapour deposition techniques [126, 127]. While maintaining a relatively long of 1.46 s and strong triplet spin polarisation, DAP:PTP exhibited an enhanced contrast of up to 40, surpassing the state of the art in NV-diamond platforms [127]. This improvement was attributed to the introduction of new low-energy non-bonding states that enhance vibrationally-induced spin-orbit coupling between the ground-state singlet and the metastable Tx state. It was also shown that direct solvent techniques can be used to grow ODMR-active DAP:PTP nanoparticles, which could be compatible with in vitro sensing experiments.
Further organic systems have demonstrated potential as FOAMs, though, to our knowledge, pulsed optically detected experiments have yet to be performed. For example, the room temperature steady-state ODMR contrast of a 1 crystal of pentacene-doped picene has been measured at 15, representing a potentially significant improvement over the PTP matrix [130]. Steady-state ODMR signals have also previously been reported at 2 K for (perdeutero)tetracene, 1,2-benzathracene, 1,2,3,4-dibenzanthracene [131] and dinaphtho-(2’,3’:1,2);(2",3":6,7)-pyrene [132]. Interestingly, work by Corvaja, Pasimeni and Giometti et al., have shown that even highly spin-dense charge-transfer (CT) co-crystals can exhibit bright room-temperature ODMR signals. Co-crystals comprised of donors such as biphenyl, fluorene, phenazine and acceptors such as 7,7’:8,8’-tetracyanoquinodimethane (TCNQ) and 1,2,3,4-tetrafluoro-TCNQ (F4TCNQ) have been studied to elucidate their triplet state dynamics [133, 134]. These materials exhibit narrow resonance lines due to intermolecular site hopping of triplet excitons. The resonances for each site can become resolved at low temperatures where hopping is not thermodynamically favoured [133]. Their high spin densities (50) are highly advantageous for quantum sensing, where the a.c. sensitivity is proportional to the , though this is often concurrent with less robust quantum spin properties compared with dilute materials such as Pc:PTP. Moreover, these materials are also promising hosts for the study of exotic spin behaviours like singlet fission and triplet-triplet annihilation [135]. For example, at room-temperature, the triplet states of Phenazine:TCNQ are predominantly formed by singlet fission and using transient nutation ESR spectroscopy the authors estimate and times of 1 s and 600 ns, respectively, even at room temperature [136].
Several ground-state radical and diradical materials with an optical readout capacity have also been demonstrated [137, 15, 16], as have materials with quintet states [138, 139, 140]. Diradical materials have the advantage of exhibiting a ground-state triplet, making their optical addressability protocol identical to NV-diamond. Optical readout is enabled by employing luminescent radical moieties based on tris(2,4,6-trichlorophenyl)methyl (TTM) [141], where the radical is sterically isolated on a carbon atom with only weak hyperfine coupling to neighbouring 13C and hydrogen atoms. As a result, these materials can benefit from relatively narrow resonance linewidths, long s, and can be embedded in frozen solutions or optically transparent PMMA foils through simple solvent-based methods [129].
Recently, even fluorescent proteins have been demonstrated as a viable spin qubit medium in vivo and in solution at room temperature. Feder et al., used ODMR spectroscopy to demonstrate that at 80 K the X-Z and Y-Z triplet transitions of enhanced yellow fluorescent protein (EYFP) demonstrated a 44 and 32 ODMR contrast, respectively [56]. This corresponds with a zero-applied field of 141 s (estimated from spin polarisation decay) and of 1.5 s. Using Carr-Purcell-Meiboom-Gill (CPMG) dynamical decoupling the effective decoherence time () reached 16 s. The authors largely circumvent overhead limitations from long triplet lifetimes (ms) using an additional near-infrared pulse to induce T1 to T2 transitions to induce reverse ISC and delayed fluorescence from S1. This interesting methodology could be suitable for other spin systems with quasi-resonant S1 and T2 electronic states, such as pentacene, to increase their sensitivity by increasing measurement repetition rates.
2.3.2 d-block molecular spin centres
Molecular FOAMs with d-block elements have only recently been demonstrated with Cr4+. The Cr4+ spin-centre was realised with several tolyl-based ligand systems where modifications to the ligand structure and corresponding ligand field strength can be used to control optical excitation frequency and the ZFS of the Cr4+ S=1 ground state [142]. To reduce intermolecular dipole coupling, the spin-active moiety is diluted in a matrix with the S=0 isostructural tin analogue in crystals grown from hexane solutions [18]. Optical spin polarisation is then achieved by exciting molecular spins from the S=1 state to the S=0 manifold, which fluorescently decays within a few microseconds to favourably populate the T± states. Initial pulsed ODMR experiments were performed on Cr4+(o-tolyl)4 spins, which benefit from an E=0 triplet and a relatively long of 0.22 ms and of 640 ns at 5 K. This permits several optical cycles to build spin polarisation and enhance the optical contrast up to 14. The coherent properties can be significantly improved using a non-isostructural Sn(4-fluoro-2-methylphenyl)4 host matrix [143]. Here, the ZFS is significantly increased, giving rise to clock transitions similar to those realised by the so-called "zero first-order Zeeman" (ZEFOZ) method, initially pioneered for lanthanide spin systems (see below). As a result, and are increased to 1.210.02 ms and 10.60.2 s at 5 K, respectively, while the pulsed contrast can be improved up to 65.
Work continues to explore alternative d-block elements where strong ligand-field interactions can enable optical excitation between spin manifolds and spin-dependent fluorescence [115, 118]. Candidate include Ni2+ [144, 145], Mo4+ [146], V3+ [147, 146]. Current challenges involve tuning the energy of the intermediate (e.g., singlet) state to provide the right conditions for spin polarisation, and identifying the appropriate ligand design to optimise photoluminescence.
2.3.3 f-block molecular spin centres
Lastly, there has been significant interest in f-block molecular systems. Here, we make a distinction between molecular systems where the ion is doped into a molecular lattice, and trapped ion systems where a single ion is levitated using electrostatic interactions to give rise to extremely long coherence times, but are not subject to qubit-quality improvements through molecular engineering [148, 149, 150].
Lanthanide spin centres benefit from highly shielded f-orbital electrons compared to d-orbital systems, leading to relatively long coherence times at low temperatures [151]. The core-like nature of f-orbitals also means, compared to d-orbital systems, that the ligand-field environment has a much weaker influence on state energies (a few THz), compared to electron-electron interactions (800 - 1600 nm) and spin-orbit coupling (10s THz) [152, 153] (see Figure 9a). Lanthanide electron spins can also exhibit strong hyperfine coupling (tens of MHz) with I=5/2 (Eu, Pr, 173Yb), I=1/2 (171Yb), I=7/2 (149Sm, 167Er) nuclei, leading to a diverse spin-optical addressibility schemes that can be further enhanced using applied magnetic fields. Moreover, strong spin-orbit coupling can lead to high magnetic anisotropy that can protect spin states from small magnetic fluctuations [154, 155]. Examples of optically active materials benefit from narrow and stable spin-dependent emission profiles at near-infrared frequencies, making them potentially compatible with conventional telecom optical fibres. Colour centres for which optical addressability has been established include Ce3+ [156], Eu3+- [157], Pr3+-[50], Er3+- [47], Yb3+- [158], and Sm3+:Y2SiO5 [159] (Figure 9c).
In the solid-state, Y2SiO5 has been favoured as host matrix due to the ability to grow large crystals with excellent optical properties using the Czochralski method [160] (Figure 9b). However, as the only naturally occurring isotope, 89Y harbours an I=1/2 nuclear spin that ultimately limits decoherence times. To reduce the impact of the spin bath, various decoupling techniques have emerged [161, 162, 45]. Perhaps the most successful is the so-called "ZEFOZ" technique, whereby a magnetic field is applied such that the magnetic-field dependence of the spin-transition frequency is very close to zero [163] (Figure 9d). In this "clock transition" regime, spins are first-order insensitive to small fluctuations in local magnetic fields. Using this technique with Eu3+:Y2SiO5, Zhong et al., demonstrated it is possible to acquire >>100 ms at 2 K, where spin-phonon coupling is negligible. Remarkably, combined with dynamic decoupling methods was measured to be 37060 mins, reaching a critical milestone whereby the distance-dependent decoherence becomes less for spin-transport than it is during light-transport of quantum states [164].
Interestingly, the larger magnetic moment of Pr3+ can give rise to "frozen core"-type behaviour, whereby local Y-spins become dephased from the bulk crystal and hence exhibit reduced dephasing influence on the Pr3+ spins [161]. Equall et al., measured homogenous field-dependent and crystal structure site-dependent linewidths between 2.5 and 0.85 kHz, and a corresponding as high as 377 s at 1.4 K [50]. Combined with the ZEFOZ method, the can reach 82 ms at 1.5 K [161] (see example Figure 9e) and with further dynamical decoupling up to 1 min can be achieved, approaching the population lifetime limit [165]. Coherence times can also be enhanced using a host matrix whereby the principal host ion (e.g., Y) has a more closely matched ionic radius to the dopant, leading to reduced crystallographic distortions. For example, in a Pr3+:La2(WO4)3 system a of 1587 ms has been measured at 4 K [51].
Spin coherence times of Er3+:Y2SiO5 have so far been shorter than the best performing Eu3+ materials with measured at 1.30.01 seconds at 1.4 K, despite exhibiting the frozen core effect [47]. However, this was achieved without using the ZEFOZ method, and therefore, these times can likely be significantly extended. Er3+ has also been studied in a Y2O3 host with reaching 140 s at 1.8 K and with an applied field of 4 T [46]. Decoherence was dominated by phonon-driven dipole-dipole interactions and the nuclear spin bath at high fields, similar to Er3+:KTiOPO4 where was measured at 200 s under similar conditions [48].
171Yb3+ holds a unique position amongst the lanthanide ions discussed so far due to its S=1/2 and I=1/2 electron spin and hyperfine structure, resulting in a simple 4-level system. Of the 171Yb3+-doped materials [53], 171Yb3+:Y2SiO5 appears to exhibit the longest spin coherence times. This material was initially studied by X-band EPR spectroscopy and presented with an electron of 5 seconds at 2.5 K with an applied field of 90 mT. quickly increases above 4 K due to Raman relaxation where becomes ultimately limited by [52]. At 2.5 K, was optimised at 1 T to 73 s and improved further by dynamical decoupling reaching 550 s. In the same experiments, the authors record nuclear and at 4.5 K of 4 and 0.35 ms, respectively. Using an optical approach, Ortu et al., improved the coherent properties of 171Yb3+:Y2SiO5 by employing the ZEFOZ method such that the electron remains above 100 s at 5.6 K and the nuclear extends to 1 ms [158]. Using a different approach, Welinski et al., demonstrated that coherence can also be extended by first polarising host nuclear spins through spin diffusion. At 2 K, the authors first excite 171Yb spins before allowing them to equilibrate over a few seconds through spectral diffusion over the inhomogeneous linewidth. The result is an effective "hole burning" in the absorption spectrum of 171Yb3+:Y2SiO5 and up to 90 nuclear spin polarisation, thereby effectively generating mK spin temperatures and improving the optical from 0.3 to 0.8 ms [168].
3 Optimising Performance of Spin-Qubit Materials
From our review, it is clear that spin-based qubit candidates demonstrate potential in several fields of quantum technology. However, unsurprisingly, their spin coherence lifetimes are limited by temperature and the spin-bath dependence of the spin properties and optical contrast between different spin states. Across all material platforms, there is significant magnetic inhomogeneity that emerges from random local spin environments. To understand the extent to which inhomogeneity infects different materials, it could be instructive to consider the ratio of / (Figure 10, the so-called "inhomogeneity parameter"). Hence, a low ratio indicates that is close to and the materials are limited by dynamic, not static, sources of decoherence. Using the available data where and were measured under similar conditions, there appears to emerge a distinct advantage for molecules and van der Waals materials at higher temperatures despite their lack of isotopic enrichment. This likely stems from the use of molecular crystals and the inherent 2D-order associated materials such as hBN, which helps to helps to ensure that all molecules "feel" the same magnetic environment. Interestingly, despite the significant difficulties associated with synthesising aligned 3D-spin defects in diamond, SnV-diamond (measured as single spins) exhibits the lowest ratio at low temperatures. Considering its robust quantum spin parameters, SnV-diamond could be a leading candidate for low-temperature applications.
Spin parameters can be improved by positional engineering of defect centres. Clustering (or the straggling) of spin-active dopants in a substrate poses spin-spin coupling from the environment (nuclear spins), which decreases . Controlled doping becomes crucial in decreasing spin density around spin-active defect centres. Eliminating unwanted spins like nuclear spins requires isotopic purity of substrate material or the careful doping of spin-active centres or defects within the host matrix. More recently, Plasma-Enhanced CVD methods have been used to achieve higher deposition rates while minimising the impact of energetic ions or electrons affecting the colour centres in NV and SnV diamonds [170]. Another novel method to precisely control the position of each defect is laser writing. Aberration-corrected optics allow for the precise positioning of vacancies in diamond systems, with a 45 success probability of a vacancy being located within 200 nm of a desired position [171].
Due to their ability to effectively engineer the placement of molecules in 3D chemical systems, materials with charge transfer or hydrogen bonding motifs may yet demonstrate advantages. Moreover, as seen with Pc:PTP, and Pr3+:Y2SiO5 vs. the La2(WO5)3 host, material engineers should avoid defect-site strain by selecting hosts with closely matched physical parameters to the defect. Finally, further improvements can be realised through dynamical decoupling methods that are tailored for each application. For example, while engineering clock transitions using magnetic fields or ZFS is appealing for quantum optics where spectral stability is prized, it is not necessarily useful for sensing or information processing. This is because the associated reduced magnetic field sensitivity would reduce the quantum operation fidelity. On the other hand, focused electromagnetic driving of parasitic impurities, such as N-centres in diamond, has yet to be significantly explored in molecular systems. Significant improvements in and would likely emerge from a combination of field-driving and pulsed refocusing methods such as CPMG.
4 Conclusion and Outlook
Optically addressable spin systems show great potential as a diverse form of qubit media for quantum sensing, communications, and information processing applications. From this review, we have attempted to benchmark the different families of materials and identify useful investigative and experimental approaches that can be translated across the field. The most significant hurdles faced by chemists and materials engineers remain the strong temperature dependence of the spin-lattice relaxation and thermal polarisation of the spin bath, which renders most d- and f-block metal systems impractical above a few kelvin, below which these materials demonstrate the longest coherence times by a significant margin. However, above liquid helium temperatures, it is clear that light-element materials such as colour centres in diamond, SiC and most recently, molecular systems demonstrate more robust quantum spin parameters. Nevertheless, significant advancements in dynamic decoupling techniques within the last 20 years and isotope engineering have enabled the realisation of remarkably competitive quantum spin properties. Within the next decade, it should be expected that further advances and knowledge transfer between investigators of different material platforms will lead to devices capable of significantly impacting society, especially in the field of magnetic field sensors and coherent quantum optics. We also hope that by collating (meta)data on the available spin qubit candidates that the field will eventually benefit from the processing and predictive power of machine learning and artificial intelligence approaches to materials discovery.
5 Acknowledgements
This work was supported by the UK Engineering and Physical Science Research Council through grants EP/V048430/1 and EP/W027542/1, and the Department of Materials, Imperial College London. The authors are also very grateful to Dr. Daan M. Arroo for fruitful discussions and advice.
6 Author contributions
CAT and MA curated the data. CAT, MO and MA contributed to the writing of the manuscript.